P=2L+2W. where L is the length and W is the width. the perimter of the problem is 24 inches. find it's dimensions if it's length is 3 inches greater than its width...thanks
Answer by mmm4444bot(95)

Hello There:
This problem tells us that we can express the length in terms of the width because we are told that the length is the same as the width plus 3 inches.
Let's assign a variable name to represent the unknown width.
w = width
Now we can write an expression for the length.
w + 3 = length
Substitute this expression for the length into the given formula for the perimeter.
P = 2*(w + 3) + 2*w
Use the distributive property to get rid of the parentheses.
P = 2*2 + 6 + 2*w
Combine the two terms containing the variable to simplify.
P = 4*w + 6
We are told that P = 24, so let's make that substitution now.
4*w + 6 = 24
Can you solve this equation for w?
Once you find the value of w, you get the length by adding 3.
~ Mark
P.S. You should get the following dimensions: 9/2 inches by 15/2 inches.